For the shape shown above, b has a value of 4

  1. Calculate the location of the x centroid
  2. Calculate the MOI of the shape about the y axis
  3. Calculate the MOI of the shape about the y axis through the x centroid

| ------- | Ix=y2/ 67 6/2 | ------- -
Answer #1

Y = x - - - > 2 - - b=4 0 F © Calculation of x-centroid x - 4 y yo, 4x = [4 - 2 y = mc L 15 o distance dx) x - y dz) x 2 of c③ MOI of the shape about y-axis Iy = 13, de 2² - J ada) 2² 73.1428 – 32 41.1428) Iy = h Moment of Inertia about y axis throug